# Liouville theorem for biharmonic functions on manifolds of nonnegative Ricci curvature

### v1

@@ Abstract.
In this paper we extend Yau’s celebrated Liouville theorem to the biharmonic case. Namely, we show that in a complete Riemannian manifold with a pole, nonnegative Ricci curvature and such that the geodesic spheres centered at the pole are mean-convex (or if the square of the distance function from the pole is convex), any biharmonic function of subquadratic growth must be harmonic, and hence, any biharmonic function of sublinear growth must be constant. Our proof relies on a new local $L^{2}$ estimate for the Laplacian of biharmonic functions combined with a mean value inequality. Examples where our theorem applies include hypersurfaces of positive sectional curvature in $\mathbb{R}^{n}$, and manifolds with a pole of nonnegative Ricci curvature whose curvature decays at infinity rapidly enough.
@@ 1. Introduction
The Liouville theorem, a landmark result of classical analysis, asserts that any bounded harmonic function on $\mathbb{R}^{n}$ is constant. In 1975, Yau [14] proved that this theorem continues to hold for complete Riemannian manifolds with nonnegative Ricci curvature, that is, every bounded harmonic function must be constant. Since then, the study of Liouville-type properties has been a central theme in geometric analysis, with deep connections to potential theory, volume growth, and curvature conditions (see Li–Schoen [8], Ni–Tam [11], Colding-Minicozzi [5], among many others).
@@
On Riemannian manifolds, of course, the situation is subtler. While the study of polyharmonic operators arises naturally in conformal geometry, most notably through the Paneitz operator, general Liouville theorems for biharmonic functions under Ricci curvature conditions appear to be largely absent from the literature. As notable exceptions we include the works Branding [2] where Liouville-type theorems for biharmonic maps between manifolds are proved under smallness conditions on different Sobolev norms. In addition, Maz’ya and Mayboroda [9, 10] obtained fundamental $L^{2}$ estimates for solutions of biharmonic Poisson problems in Euclidean domains, which inspired our search for analogous inequalities in the Riemannian setting. Also, previous work by the authors [4] extends Milnor’s criterion for parabolicity of surfaces to biharmonic functions, and a curvature regime where any bounded biharmonic function must be harmonic is identified. We believe that the study of Liouville-type properties and its relation with geometric invariants for higher order elliptic operators is a fundamental question in Geometric Analysis.
@@
is a diffeomorphism. We shall also include another geometric hypothesis. Either
  • (H1)
    the spheres around $p$ are mean-convex with respect to the radial vector field, or
  • (H2)
    The square of the distance function from $p$, that we denote by $d$, is convex.
Then we have the following
@@
Let $(M,g)$ be a complete Riemannian manifold with $\mbox{Ric}\geq 0$. Assume that either (H1) or (H2) above holds.Then any biharmonic function $u\in C^{4}(M)$ of subquadratic growth must be harmonic. In particular, any biharmonic function of sublinear growth must be constant.
Assumptions (H1) and (H2) are actually natural geometric hypotheses. If $M$ is a complete Riemannian manifold of positive sectional curvature diffeomorphic to $\mathbb{R}^{n}$, then, by the soul theorem, it has a pole and the spheres around that pole are mean convex, so assumption (H1) holds: this shows, for instance, that surfaces such as elliptic paraboloids, and a sheet of a two sheeted hyperboloid satisfy assumption (H1). On the other hand, assumption (H2) has been considered by Kasue in [7], where he shows that if the manifold has a pole and the curvature at infinity decays fast enough, then (H2) holds. We do not know whether (H1) and (H2) can be removed from the hypotheses of Theorem 1: We conjecture that it can be done.
@@ 2.1. Some Technical Lemmas
We start with a classical cut-off construction, which is the key in the proof of our main theorem. Here the geometric assumptions (H1) or (H2) become fundamental.
@@
Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$. Assume that assumption (H1) or (H2) holds. Let $0<r<R$. Set $B_{\rho}:=B_{\rho}(p)$. Assume $\mbox{Ric}_{g}\geq-(n-1)K\,g$ on $B_{R}$ for some $K\geq 0$. Then there exists $\chi\in C_{c}^{\infty}(B_{R})$ such that
@@
By the Laplacian comparison under the curvature inequality $\mbox{Ric}\geq-(n-1)K$, $\Delta d\leq(n-1)f_{K}(d),$ with the convention that $f_{K}(d)=\sqrt{K}\coth(\sqrt{K}\,d)$ if $K>0$ and $f_{K}(d)=1/d$ if $K=0$. Thus, assumption (H1) implies that
@@
Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$. Assume that assumption (H1) or (H2) holds. Set $B_{\rho}:=B_{\rho}(p)$, assume $\mbox{Ric}_{g}\geq 0$ on $B_{R}$, and let $0<r<R$. Then there exists $\varphi\in C_{c}^{\infty}(B_{r})$ such that $0\leq\varphi\leq 1$, $\varphi=1$ on $B_{s}$, and for $\dfrac{1}{8}r<s<r<R$,
@@ 2.2. Local energy estimates and Cacciopoli-type inequalities
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with a pole $p$ with $\mathrm{Ric}\geq 0$ and such that (H1) or (H2) holds. Let $B_{\rho}\colon=B_{\rho}\left(p\right)$. Fix $a>0$ and let $u\in C^{\infty}(B_{2a})$ be biharmonic, i.e. $\Delta^{2}u=0$ on $B_{2a}$. Then there exists a constant $C=C(n)>0$ such that, for every pair of radii $0<\dfrac{1}{8}r<s<r\leq 2a$,
@@
Let $(M,g)$ be a smooth Riemannian manifold with a pole $p\in M$, such that $\mbox{Ric}\geq 0$, and such that (H1) or (H2) holds. Let $0<\dfrac{1}{8}r<s<r<\infty$ and let $B_{\rho}:=B_{\rho}(p)$. If $u\in C^{\infty}(B_{2r})$ is biharmonic, i.e. $\Delta^{2}u=0$ in $B_{2r}$, then there exists a constant $C=C(n)>0$ such that

### v3

@@ Abstract.
In this paper we extend Yau’s celebrated Liouville theorem to the biharmonic case. Namely, we show that in a complete Riemannian manifold with a pole and nonnegative Ricci curvature, any biharmonic function of subquadratic growth must be harmonic, and hence, any biharmonic function of sublinear growth must be constant. Our proof relies on a new local $L^{2}$ estimate for the Laplacian of biharmonic functions combined with a mean value inequality. Examples where our theorem applies include hypersurfaces of positive sectional curvature in $\mathbb{R}^{n}$, and manifolds with a pole of nonnegative Ricci curvature whose curvature decays at infinity rapidly enough.
@@ 1. Introduction
The Liouville theorem, a landmark result of classical analysis, asserts that any bounded harmonic function on $\mathbb{R}^{n}$ is constant. In 1975, Yau [13] proved that this theorem continues to hold for complete Riemannian manifolds with nonnegative Ricci curvature, that is, every bounded harmonic function must be constant. Since then, the study of Liouville-type properties has been a central theme in geometric analysis, with deep connections to potential theory, volume growth, and curvature conditions (see Li–Schoen [8], Li–Tam [9], Colding-Minicozzi [5], among many others).
@@
On Riemannian manifolds, of course, the situation is subtler. While the study of polyharmonic operators arises naturally in conformal geometry, most notably through the Paneitz operator, general Liouville theorems for biharmonic functions under Ricci curvature conditions appear to be largely absent from the literature. As notable exceptions we include the works Branding [2] where Liouville-type theorems for biharmonic maps between manifolds are proved under smallness conditions on different Sobolev norms. In addition, Maz’ya and Mayboroda [10] obtained fundamental $L^{2}$ estimates for solutions of biharmonic Poisson problems in Euclidean domains, which inspired our search for analogous inequalities in the Riemannian setting. Also, previous work by the authors [4] extends Milnor’s criterion for parabolicity of surfaces to biharmonic functions, and a curvature regime where any bounded biharmonic function must be harmonic is identified. We believe that the study of Liouville-type properties and its relation with geometric invariants for higher order elliptic operators is a fundamental question in Geometric Analysis.
@@
is a diffeomorphism. Then we have the following.
@@
Let $(M,g)$ be a complete Riemannian manifold with $\mbox{Ric}\geq 0$. Then any biharmonic function $u\in C^{4}(M)$ of subquadratic growth must be harmonic. In particular, any biharmonic function of sublinear growth must be constant.
@@ 2.1. Some Technical Lemmas
We start with a classical cut-off construction, which is the key in the proof of our main theorem. But first, we need to make an observation. If $(M,g)$ has a pole and $\mbox{Ric}_{g}\geq 0$ then $\Delta d\geq 0$, where $r$ is the distance function from the pole. The reason is that, along a geodesic $\gamma:\left[0,\infty\right)\longrightarrow M$ starting at the pole, the Laplacian of the distance function from the pole $S(t)=(\Delta d)(\gamma(t))$ satisfies the Ricatti inequality
$S^{\prime}\leq-\frac{S^{2}}{n-1}-\mbox{Ric}_{g}(\dot{\gamma},\dot{\gamma}),$
and hence
$S^{\prime}\leq-\frac{S^{2}}{n-1},$
so that if $S$ becomes negative at some point, then there would be a $t_{0}>0$ such that
$\lim_{t\rightarrow t_{0}^{+}}S(t)=-\infty,$
contradicting that the geodesic starts at a pole.
@@
Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$. Let $r$ be such that $0<r<R$. Set $B_{\rho}:=B_{\rho}(p)$. Assume $\mbox{Ric}_{g}\geq-(n-1)K\,g$ on $B_{R}$ for some $K\geq 0$. Then there exists $\chi\in C_{c}^{\infty}(B_{R})$ such that
@@
By the Laplacian comparison under the curvature inequality $\mbox{Ric}\geq-(n-1)K$, $\Delta d\leq(n-1)f_{K}(d),$ with the convention that $f_{K}(d)=\sqrt{K}\coth(\sqrt{K}\,d)$ if $K>0$ and $f_{K}(d)=1/d$ if $K=0$. Thus,
@@
Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$. Set $B_{\rho}:=B_{\rho}(p)$, assume $\mbox{Ric}_{g}\geq 0$ on $B_{R}$, and let $0<r<R$. Then there exists $\varphi\in C_{c}^{\infty}(B_{r})$ such that $0\leq\varphi\leq 1$, $\varphi=1$ on $B_{s}$, and for $\dfrac{1}{8}r<s<r<R$,
@@ 2.2. Local energy estimates and Cacciopoli-type inequalities
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with a pole $p$ with $\mathrm{Ric}\geq 0$. Let $B_{\rho}\colon=B_{\rho}\left(p\right)$. Fix $a>0$ and let $u\in C^{\infty}(B_{2a})$ be biharmonic, i.e. $\Delta^{2}u=0$ on $B_{2a}$. Then there exists a constant $C=C(n)>0$ such that, for every pair of radii $0<\dfrac{1}{8}r<s<r\leq 2a$,
@@
Let $(M,g)$ be a smooth Riemannian manifold with a pole $p\in M$, such that $\mbox{Ric}\geq 0$. Let $0<\dfrac{1}{8}r<s<r<\infty$ and let $B_{\rho}:=B_{\rho}(p)$. If $u\in C^{\infty}(B_{2r})$ is biharmonic, i.e. $\Delta^{2}u=0$ in $B_{2r}$, then there exists a constant $C=C(n)>0$ such that